# An Adverse Weather-Immune Scheme with Unfolded Regularization and Foundation Model Knowledge Distillation for Street Scene Understanding

### v1

@@ IV-A3 Implementation Details
Our deep learning model was implemented using the Pytorch framework, with experiments conducted on two NVIDIA GeForce 4090 GPUs. For the architecture, we implement the proposed AdvImmu based on ResNet [35] and ASSP perception head [32]. We select the Adam optimizer for training, configuring it with Betas values of 0.9 and 0.999, and set the weight decay at 1e-4. The training was executed with a batch size of 8 and a learning rate of 3e-4.
IV-B AdvImmu’s Overall Performance Evaluation
@@
Tab. I compares the performance of the AdvImmu against DeepLabv3+ [32], BiSeNetV2 [33] and SegNet [34]. Across all metrics, AdvImmu generally outperforms DeepLabv3+, BiSeNetV2 and SegNet on both datasets. For example, on Apolloscapes dataset, AdvImmue outperforms DeepLabv3+, BiSeNetV2 and SegNet by (59.35 - 26.58) / 26.58 = 123.29%, (59.35 - 22.92) / 22.92 = 158.94% and (59.35 - 21.01) / 21.01 = 182.48% in mIoU, respectively; on CARLA_ADV dataset, AdvImmue outperforms DeepLabv3+, BiSeNetV2 and SegNet by (69.58 - 36.90) / 36.90 = 88.56%, (69.58 - 28.80) / 28.80 = 141.60% and (69.58 - 31.67) / 31.67 = 119.70% in mIoU, respectively. This suggests that AdvImmu should be more effective compared to the SOTA DeepLabv3+ model.
LABEL:Fig.perf_comp_inf compares the inference performance of AdvImmu against DeepLabv3+ (performing best in three baselines) in real driving test. Notably, AdvImmu consistently demonstrates better performance relative to DeepLabv3+, which can be supported by following aspects: (I) AdvImmu consistently achieves higher scores especially in mF1 and mIoU across all sequential frames, indicating better accuracy. (II) AdvImmu shows small performance variance across varied test conditions, such as fog, cloudy, rainy, dark, and combinations thereof, suggesting robustness against adverse weather conditions.
TABLE II: The effect of LSM depth on AdvImmu performance
@@
IV-C In-depth Insights of AdvImmu
IV-C1 LSM Depth on Performance
Tab. II explores the effect of the LSM depth on performance. Across different LSM depths (Depth-1 through Depth-4), the performance metrics show certain variation, indicating that changes in depth have a little impact on AdvImmu’s performance. LABEL:Fig.LSM_depth_comp corroborates this finding, showing that all depths reach different performance levels, with slight fluctuations. However, the subtle variations in performance across different depths indicate that both too shallow (Depth-1) and too deep (Depth-4) configurations do not consistently yield the best results. Depths-3 generally show more stable and slightly improved performances, suggesting that moderate depths could be optimal. This implies that an intermediate LSM depth could potentially offer a better compromise between computation complexity and performance.
@@ IV-C2 EF vs LF
LABEL:Fig.EF_vs_LF presents the comparison between EF and LF. LABEL:Fig.EF_arch and LABEL:Fig.LF_arch illustrate the architecture of EF and LF of AdvImmu. LABEL:Fig.EF_vs_LF_perf compares the performance of EF and LF across four metrics, where LF performs slightly better in mIoU and mPre, while EF has a slight advantage in mRec and mF1. Overall, the differences between the two fusion methods are minimal, suggesting that both approaches offer comparable performance with slight variations depending on the specific metric being evaluated.
@@ IV-C3 Enabling GSM vs Disabling GSM
Tab. III demonstrates that enabling GSM significantly improves the model performance across all measured metrics. LABEL:Fig.GSM_comp shows that with GSM enabled, the performance metrics are consistently higher and more stable across 200 epochs, compared to the lower and more fluctuating metrics when GSM is disabled. Overall, enabling GSM leads to better and more reliable performance.
TABLE III: The effect of GSM on AdvImmu performance
@@
TABLE IV: The regularizer comparison
@@ IV-C4 Vanilla regurlarizers (VRs) vs URs
Tab. IV and LABEL:Fig.regularizer_comp compare the performance of the following cases: (I) only cross entropy loss (denoted as CE); (II) cross entropy plus vanilla regularizers (denoted as VRs); (III) cross entropy plus 2-layer unfolded regularizers (denoted as URs ($L=2$)); (IV) cross entropy plus 5-layer unfolded regularizers (denoted as URs ($L=5$)). By comparing these cases, we can observe the following patterns: (I) Tab. IV shows a consistent improvement in performance when additional regularization techniques are added to the basic CE loss, leading to better scores in mIoU, mPre, mRec, and mF1. (II) Tab. IV also demonstrates that URs outperform corresponding VRs in almost all metrics. For example, URs ($L=2$) and URs ($L=5$) achieve mIoUs of 52.13 and 59.35, respectively, compared to 31.49 achieved by VRs. (III) Tab. IV additionally shows that the more layers the regularizers are unfolded, the better the performance obtained. For instance, URs ($L=5$) achieve an mIoU of 59.35, better than the 52.13 achieved by URs ($L=2$). These patterns can be visually confirmed in LABEL:Fig.regularizer_comp. In addition, LABEL:Fig.regularizer_comp illustrates that URs ($L=5$) converges faster than URs ($L=2$).
@@ IV-D SBICAC Evaluation
LABEL:Fig.cluster_comp presents a qualitative comparison between SBICAC and other clustering algorithms. To quantitatively evaluate the proposed SBICAC against such baselines, we compare the mIoU between the ground truth and the respective masks generated by such baselines. However, to overcome the issue of the mismatch of cluster IDs, an approach is structured as follows:
@@
LABEL:Fig.clutering_perf_mIoU shows a quantitative comparison of four clustering algorithms. The proposed SBICAC algorithm performs the best with an mIoU of 57.00%. It is followed by SpectralClustering at 53.10%, AgglomerativeClustering at 46.80%, and KMeans with the lowest mIoU of 30.40%. This demonstrates that SBICAC outperforms the other three algorithms. On the other hand, we utilize both the ground truth and SBICAC-generated pseudo-labels to train the proposed AdvImmu. As shown in LABEL:Fig.clutering_perf_training, the validation results in mIoU reveal that training with SBICAC-generated labels can achieve performance comparable to, or even surpassing, that of the ground truth. This superior performance is attributed to the fact that the generated labels are coarser than the ground truth (i.e., easier to classify).
@@ V Conclusion
AD perception in various adverse weather conditions is challenging. This work proposed AdvImmu to enhance the model generalization across multiple adverse weather conditions by considering local temporal correlation. It processed frame sequences using InsU, IntU and DU to capture key information, consistent background and dynamic changes, and then shuffles these processed features to avoid overfitting and improve generalization. The training of AdvImmu requires frame annotations, not usually available in AD. To overcome this, we integrated the SAM and SBICAC clustering algorithm to support training. Our tests show that AdvImmu outperforms existing methods significantly.

### v3

@@ IV-A3 Implementation Details
The proposed AdvImmu model is implemented using the Pytorch framework, with experiments conducted on two NVIDIA GeForce 4090 GPUs. For the architecture, we implement the proposed AdvImmu based on ResNet [77] and ASSP perception head [78]. We compare the proposed AdvImmu with other 8 baselines, i.e., DeepLabv3+ [78], BiSeNetV2 [13], SegNet [79], TopFormer [80], SeaFormer [81], BASeg [82], HRDA [15], and AttaNet [83]. All these baselines are also implemented using Pytorch. We train the proposed AdvImmu and such baselines based on aforementioned datasets from scratch. For the optimization, we select the Adam optimizer for training, configuring it with Betas values of 0.9 and 0.999 and setting the weight decay to 1e-4. The training is executed with a batch size of 8 and a learning rate of 3e-4. For the InfoNCE regularizer, the hyperparameter $\tau$ is set to 0.5. For URs, the number of unfolded layers $K$ is set to 5.
IV-B AdvImmu’s Overall Evaluation
In this section, we will conduct extensive expriments to explore the proposed AdvImmu from following aspects: performance, convergence, compleixty (measured by Floating Point Operations (FLOPs)), and inference delay (represented by Frames Per Second (FPS)).
IV-B1 AdvImmu’s Performance
@@
Table I compares the performance of the AdvImmu against DeepLabv3+, BiSeNetV2, SegNet, AttaNet, BASeg, HRDA, SeaFormer and TopFormer, across Apolloscapes, CARLA_ADV, Shift, SynthiaSF, Cityscapes and GTA5 datasets quantitatively. Notably, in this table, bold font denotes the best performance model and underline denotes the second best performance model with respect to evaluation metrics. From Table I, we can observe that for almost all datasets, AdvImmu generally outperforms all other baselines across almost all metrics. Specifically, taking Apolloscapes dataset as example, AdvImmue outperforms DeepLabv3+, BiSeNetV2, SegNet, AttaNet, BASeg, HRDA, SeaFormer and TopFormer by (59.35 - 26.58) / 26.58 = 123.29%, (59.35 - 22.92) / 22.92 = 158.94%, (59.35 - 21.01) / 21.01 = 182.48%, (59.35 - 20.59) / 20.59 = 188.25%, (59.35 - 20.14) / 20.14 = 194.69%, (59.35 - 21.55) / 21.55 = 175.41%, (59.35 - 20.34) / 20.34 = 191.79%, and (59.35 - 20.41) / 20.41 = 190.79% in mIoU, respectively. This suggests that AdvImmu showcases significant performance improvement compared to other SOTA models.
Table II compares the inference performance of the proposed AdvImmu against other models qualitatively under various adverse weather conditions, such rainy, dark, foggy, cloudy, etc. From Table II, we can observe following patterns: (I) The proposed AdvImmu achieves the best performance across various adverse weather conditions and its prediction can approximately approach to the ground truth. For example, under the rainy condition (the third row), AdvImmu’s predition is quite similar to the ground truth, while others’ prediction have various weaknesses, such as class confusion (e.g., HRDA), boundary uncertainty (e.g., DeepLabv3+), small object errors (e.g., BASeg), etc. (II) Some models (e.g., HRDA) can predict well in clear weather (e.g., the first row), however, suffers from a significant performance decline under adverse conditions (e.g., the third row).
From both Table I and Table II, we can find that BASeg achieves relatively competitive performance to the proposed AdvImmu and shows overwhelming advantages over other baselines. This is attributed to the consideration of extra boundary information in the training of BASeg, which can be illustrated in LABEL:Fig.baseg_boundary. LABEL:Fig.baseg_2774 indicates an exmaple of image and LABEL:Fig.baseg_2774_bounds denotes the corresponding boundary information. Obviously, these boundary information can assist BASeg model to overcome weaknesses, such as boundary uncertainty, class confision, etc. Thus, BASeg can achieve competitive performance underer various weather conditions. However, incorporation of extra boundary information is a double-edged sword and leads to obvious prediction noise along boundaries, which is demonstrated in LABEL:Fig.baseg_2774_prediction and the third column of Table II.
IV-B2 AdvImmu’s Convergence
LABEL:Fig.Convergence_comp compares the convergence of the proposed AdvImmu against all other baselines. From LABEL:Fig.Convergence_comp, we can conclude following patterns: (I) Across almost all metrics, the proposed AdvImmu converges fastest compared to all the other baselines. For instance, taking mIoU as example, AdvImmu reaches to the maximum mIoU at around the 18-th epochs whereas other baselines reach their maximum value later than the 18-th epochs. (II) DL network architecture imposes crucial impact on convergence. Specifically, CNN-based models (e.g., AdvImmu, BASeg, DeepLabv3+, etc.) generally convergence faster than Transformer-based models (e.g., HRDA, SeaFormer, TopFormer, etc.).
IV-B3 AdvImmu’s Complexity and Inference Delay
Table III compares the FLOPs (using GFLOPs instead for being straightforward) and FPS of AdvImmu against other baselines. We can figure out following insights: (I) Models like TopFormer and BiSeNetV2 achieve high FPS with lower computational demands, making them suitable for real-time applications. (II) HRDA and BASeg have the highest GFLOPs (823.57 and 562.31, respectively) but suffer from low FPS, indicating that they are computationally intensive and slower. (III) AdvImmu has a GFLOPs similar to DeepLabv3+ but a significantly lower FPS (28.11), which is caused by AdvImmu’s loading heavy input and unfolding regularizers.
IV-C In-depth Insights of AdvImmu
In this section, we will conduct comprehensive experiments and empirical analyses to investigate how LSM, GSM, URs, and feature fusion type impact the performance of the proposed AdvImmu.
TABLE III: The comparison of GFLOPs and FPS of AdvImmu against other competitors
       Methods        GFLOPs        FPS
       DeepLabv3+        50.58        297.06
       BiSeNetV2        36.46        545.84
       SegNet        327.93        65.54
       AttaNet        23.91        506.40
       BASeg        562.31        35.60
       HRDA        823.57        9.69
       SeaFormer        13.77        234.00
       TopFormer        3.56        490.38
       AdvImmu (Ours)        50.02        28.11
TABLE IV: The effect of LSM depth on AdvImmu performance
@@ IV-C In-depth Insights of AdvImmu
IV-C1 LSM Depth on AdvImmu Performance
LSM depth represents the number of consecutive frames prior to current frame. It can impose effect on IntU and DU, and further affects LSM and AdvImmu. In this experiment, we compare four cases with setting LSM depth to 1, 2, 3 and 4, and denote them as Depth-1, Depth-2, Depth-3, and Depth-4, respectively. Table IV compares the effect of the LSM depth on AdvImmu performance. From Table IV, we can find that across different LSM depths (Depth-1 through Depth-4), the performance metrics show certain variation, indicating that changes in LSM depth have impact on AdvImmu’s performance. LABEL:Fig.LSM_depth_comp corroborates this finding visually, showing that all depths reach different performance levels with slight fluctuations. However, the subtle variations in performance across different LSM depths indicate that neither too shallow (Depth-1) nor too deep (Depth-4) configurations do not consistently yield the best results. Depth-3 generally shows more stable and slightly improved performances, suggesting that moderate depths could be optimal. This implies that an intermediate LSM depth could potentially offer a better compromise between computation complexity and performance.
@@ IV-C2 EF vs LF
EF and LF are two policies for fusing features from InsU, IntU, and DU. They both have respective pros and cons. Specifically, EF generally combines features at an early stage, allowing the model to learn interactions between InsU, IntU, and DU from the start. However, EF maybe suffers from overfitting due to the high dimensionality of combined features. In contrast, LF can preserve unique features of InsU, IntU, and DU, allowing specialized processing, whereas LF involves intensive computation and has less opportunity for learning interactions. In order to investigate how EF and LF affect the AdvImmu performance, we conduct experiments to compare their performance. LABEL:Fig.EF_vs_LF presents the comparison between EF and LF. Specifically, LABEL:Fig.EF_arch and LABEL:Fig.LF_arch illustrate the model architecture of EF and LF of AdvImmu, respectively. LABEL:Fig.EF_vs_LF_perf compares the inference performance of EF and LF across four metrics, where LF performs better than EF in all adopted metrics. Overall, although LF generally outperforms EF, LF has more learnable parameters and thus requires more powerful computational device and more training time. Therefore, we should take both model performance and training cost into account to trade off them when we choose which one actually best fits our needs.
@@ IV-C3 Enabling GSM vs Disabling GSM
GSM is designated to shuffle LSM segments to prevent the overfitting to specific temporal patterns. This is a key element to improve the across-weather performance of AdvImmu under various adverse weather conditions. To figure out how much GSM influences the performance of the proposed AdvImmu, we conduct experiments to compare the performance of enabling and disabling GSM. Table V demonstrates that enabling GSM significantly improves AdvImmu’s inference performance across all measured metrics. Specifically, we can consolidate the performance improvement from following two aspects: (I) Compared to disabling GSM, enabling GSM enhances all performance metrics with a large margin. Taking the mIoU as example, the performance improvement is (41.52 - 27.12) / 27.12 = 53.10%. (II) Compared to disabling GSM, enabling GSM improves the performance stability across multiple adverse weather conditions. Still taking mIoU as example, the standard variance reduces with the ratio of (1.84 - 0.77) / 1.84 = 58.15%. Such findings can further be confirmed visually by LABEL:Fig.GSM_comp. Concretely, LABEL:Fig.GSM_comp shows that with GSM enabled, the performance metrics are consistently higher and more stable across 200 epochs, compared to the lower and more fluctuating metrics when GSM is disabled. Overall, enabling GSM leads to better and more reliable performance.
TABLE V: The effect of GSM on AdvImmu performance
@@
TABLE VI: The regularizer comparison
@@ IV-C4 Vanilla regurlarizers (VRs) vs URs
Table VI and LABEL:Fig.regularizer_comp compare the performance of the following cases: (I) only cross entropy loss (denoted as CE); (II) cross entropy plus vanilla regularizers (denoted as VRs); (III) cross entropy plus 2-layer unfolded regularizers (denoted as URs ($L=2$)); (IV) cross entropy plus 5-layer unfolded regularizers (denoted as URs ($L=5$)). By comparing these four cases, we can observe the following patterns: (I) Table VI shows a consistent improvement in performance when the proposed image-level regularizer and inter-class contrastive regularizers are added to the basic CE loss, leading to better scores in mIoU, mPre, mRec, and mF1. (II) Table VI also demonstrates that URs outperform corresponding VRs in almost all metrics. For example, URs ($L=2$) and URs ($L=5$) achieve mIoUs of 52.13 and 59.35, respectively, compared to 31.49 achieved by VRs. (III) Table VI additionally shows that the more layers the regularizers are unfolded, the better the performance obtained. For instance, URs ($L=5$) achieves an mIoU of 59.35, better than the 52.13 achieved by URs ($L=2$). These patterns can be visually confirmed by LABEL:Fig.regularizer_comp. In addition, LABEL:Fig.regularizer_comp illustrates that URs ($L=5$) converges faster than URs ($L=2$).
TABLE VII: Output comparison among CE, VRs, URs (L=2) and URs (L=5) for specific channels
CE VRs URs (L=2) URs (L=5)
Chan1(Car)
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Chan13(Fence)
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Subsequently, in order to explore how URs improve the prediction performance, we compare the output of above four cases in a channel-wise manner. LABEL:Fig.chan_raw and LABEL:Fig.chan_gt denote an example image and the corresponding ground truth, respectively. Section IV-C4 showcases the channel-wise output of above four cases, taking the example image as input. Notably, Section IV-C4 only exemplifies the first channel (i.e., Car) and 13-th channel (i.e., Fence) out of 23 channels. From Section IV-C4, we can find that: (I) For both channels, pure CE can generally supervise the model to learn the most features of Car and Fence, but still mix other channels’ features (i.e., noises) with current channel features. For example, the light blue areas actually represent features of other channels and are mixed up together. (II) VRs can effectively draw features of the same class closer while push features of different classes further apart. Specifically, for both channels, the features of Car (or Fence) become more contrastive relative to features of remaining parts. However, VRs’ supervised features still exist noises. (III) URs enhance the learned features significantly. Precisely, the features of the Car (or Fence) are nearly identical, while those of the other areas are almost uniform, making the contrast between these two parts much clearer. Notably, as the number of the unfolded layers of URs increases, this trend becomes more pronounced, which can be demonstrated by comparing the outputs of URs (L = 2) and URs (L = 5).
Figure 23: mIoU between ground truth and various clustering algorithm-generated pseudo-labels.
@@ IV-D SBICAC Evaluation
LABEL:Fig.cluster_comp presents a qualitative comparison of generated pseudo-labels between the proposed SBICAC and other off-the-shelf clustering algorithms. To quantitatively evaluate the proposed SBICAC against such baselines, we compare the mIoU between the ground truth and the respective masks generated by such baselines. However, in general, the semantic class IDs within the generated masks by those clustering algorithms are not consistent with those within the ground truth. To overcome the mismatch of semantic class IDs for measuring the aforementioned mIoU, an approach is structured as follows:
@@
Fig. 23 shows a quantitative comparison of four clustering algorithms. The proposed SBICAC algorithm performs the best with an mIoU of 57.00%. It is followed by SpectralClustering at 53.10%, AgglomerativeClustering at 46.80%, and KMeans with the lowest mIoU of 30.40%. This demonstrates that SBICAC outperforms the other three clustering algorithms.
On the other hand, we utilize both the ground truth and SBICAC-generated pseudo-labels to train the proposed AdvImmu. Table VIII compares the inference performance of models trained by using ground truth and SBICAC-generated pseudo-labels, respectively. We can find that the models trained with SBICAC-generated labels can achieve performance comparable to, or even surpassing, the models trained with the ground truth. This superior performance is attributed to the fact that the generated labels are coarser than the ground truth (i.e., easier to classify). Thus, we can conclude that the proposed SAM+SBICAC can effectively generate pseudo-labels for the proposed AdvImmu method.
TABLE VIII: The performance comparison of models trained using ground truth and SBICAC-generated pseudo-labels
Method mIoU mPre mRec mF1
Ground Truth SAM+SBICAC Ground Truth SAM+SBICAC Ground Truth SAM+SBICAC Ground Truth SAM+SBICAC
DeepLabv3+ 26.58$\pm$0.49 24.62$\pm$0.06 30.23$\pm$0.89 26.88$\pm$0.05 31.14$\pm$0.57 32.17$\pm$0.13 32.32$\pm$0.59 28.98$\pm$0.07
BiSeNetV2 22.92$\pm$0.86 23.02$\pm$0.25 27.85$\pm$1.20 24.93$\pm$0.19 27.10$\pm$1.11 30.20$\pm$0.36 27.12$\pm$1.57 26.95$\pm$0.26
SegNet 21.01$\pm$0.51 24.51$\pm$1.15 24.80$\pm$1.11 27.89$\pm$1.71 24.96$\pm$0.54 31.69$\pm$1.22 24.60$\pm$0.70 29.23$\pm$1.48
AttaNet 20.59$\pm$0.18 25.09$\pm$0.23 25.88$\pm$0.31 28.66$\pm$0.19 25.38$\pm$0.22 32.46$\pm$0.27 25.55$\pm$0.19 29.99$\pm$0.24
BASeg 20.14$\pm$0.18 32.10$\pm$0.15 31.72$\pm$0.29 43.05$\pm$0.42 25.26$\pm$0.20 39.46$\pm$0.21 25.75$\pm$0.31 40.33$\pm$0.16
HRDA 21.55$\pm$0.27 16.86$\pm$0.69 34.13$\pm$0.31 29.40$\pm$1.40 26.09$\pm$0.34 23.12$\pm$0.66 26.40$\pm$0.31 21.88$\pm$0.70
SeaFormer 20.34$\pm$0.16 22.04$\pm$0.12 25.33$\pm$0.21 23.85$\pm$0.09 25.03$\pm$0.13 28.91$\pm$0.22 24.26$\pm$0.16 25.84$\pm$0.11
TopFormer 20.41$\pm$0.00 23.03$\pm$0.30 25.19$\pm$0.45 24.86$\pm$0.40 25.20$\pm$0.22 30.33$\pm$0.40 24.28$\pm$0.20 26.95$\pm$0.44
AdvImmu 59.35$\pm$0.57 63.97$\pm$2.20 81.38$\pm$2.22 86.60$\pm$0.99 60.76$\pm$0.49 65.63$\pm$1.77 65.28$\pm$0.55 70.78$\pm$1.57
@@ V Conclusion
AD street scene semantic understanding in various adverse weather conditions is challenging. Widely used domain adaption solution suffered from either being dependent on tough-to-collect reference images or significant performance drop in the scenario of mixup of multiple adverse weather conditions. To tackle such weaknesses of domain adaption, this work proposed AdvImmu to enhance the model generalization across multiple adverse weather conditions by considering local temporal correlation. It processes frame sequences using InsU, IntU and DU to capture key information, consistent background and dynamic changes, and then shuffles these processed features to avoid overfitting to specific temporal patterns and to improve generalization. To overcome the difficult-to-obtain frame annotations, we integrated the SAM and SBICAC clustering algorithm to support training. Extensive experiments showed that AdvImmu outperforms existing methods significantly in adopted metrics. The future works include following aspects: (I) We plan to extent the research to other tasks in the AD scenario, such as object detection, and end-to-end driving. (II) We also plan to extend the research to multi-vehicle system by incorporating federated learning paradigm.